OpenStudy (anonymous):

(4-i)^2 Simplify the expression please help

OpenStudy (anonymous):

(4-i)(4-i) =16-2i+i^2 we know that by definition i=sqrt(-1)

OpenStudy (anonymous):

16-2i+i^2 =16-2i-1 =15-2i